If a resistor is to carry 2 A of current and handle 100 W of power , how many ohms must it be ..? Assume that the voltage can be adjusted to any required value.


1
Expert's answer
2021-04-08T09:01:49-0400

Since current of resistor is given as, "I =2A"

Power, "P=100W"

Resistance of the resistor, "R = \frac{P}{I^2}= \frac{100}{4} = 25\Omega"